# PRIMA.CPP: Speeding Up 70B-Scale LLM Inference on Low-Resource Everyday Home Clusters

### v1

@@ 4.2 Ablation Study on Prefetching, Halda and Piped-ring Parallelism
Llama 3-8B 2.0% 20.0% 51.3% 42.5% 13.5% 12.8% 55.8% 12.8% 5.3% 5.4% 2.7% $\leq$1.0%
Llama 3-14B 2.5% - - - - - - - 5.3% 4.3% 2.2% $\leq$1.0%
Llama 1-30B 8.0% - - - - - - - 3.0% 5.7% 2.9% $\leq$1.0%
Llama 3-45B 3.9% - - - - - - - 4.9% $\leq$1.06.0% $\leq$1.0%
Llama 3-60B 5.5% - - - - - - - 6.3% 4.7% 4.7% $\leq$1.0%
Llama 1-65B 15.6% - - - - - - - 3.9% $\leq$1.0$\leq$1.0$\leq$1.0%
Llama 3-70B 6.0% OOM OOM OOM OOM OOM OOM OOM 4.7% 4.8% 4.8% $\leq$1.0%
4.3 Low Memory Pressure for Better Experience
Memory pressure is crucial for user experience because high pressure can slow down apps or even crash the device. Imagine if an LLM app will cause home devices to freeze, would you use it? Clearly, no. However, this is ignored by existing systems. Appendix A.6 shows devices’ memory footprint to illustrate the workload distribution and explain why prima.cpp is faster. However, a higher value does not indicate higher pressure, as reclaimable memory like page cache is included but can be freed instantly by the OS. To better measure memory pressure, we define it as the reduction in mem_available during runtime relative to mem_total. For example, 2 GiB decrease in mem_available on an 8 GiB device results in 25% pressure. As mem_available includes free and reclaimable pages, its reduction contains only non-reclaimable pages, so it is memory pressure.
Table 4 shows the memory pressure caused by llama.cpp, exo, dllama, and prima.cpp on each device. Prima.cpp and llama.cpp have low memory pressure, using only a small amount of mem_used for key-value cache and compute buffer. Since they load model weights via mmap, the majority of resident memory is held in the page cache and can be instantly released by the OS without affecting other apps. In contrast, exo and dllama keep model weights in mem_used, causing high memory pressure. This forces the OS to free inactive pages, compress memory, swap data to disk, potentially slowing down apps, leading to system lag, or OOM for larger models. Overall, exo and dllama prioritize resources for the LLM app at the cost of other apps; prima.cpp and llama.cpp prioritize user experience, keeping low memory pressure and thus better suited for user devices. Therefore, prima.cpp is the best choice for running larger models on home devices ($>$30B in our case).
@@ 5 Conclusion
This paper introduces prima.cpp, a distributed inference system to run 70B-scale LLMs on everyday home devices. The core design is a piped-ring parallelism with prefetching and a scheduler to split model layers to heterogeneous devices. Prima.cpp uses mmap to manage model weights, preventing OOM for any model size but introducing disk latency. To hide this latency, prima.cpp employs prefetching and piped-ring parallelism, which overlaps disk loading and avoiding the “prefetch-release" effect. To further speed up inference, we mathematically model token latency by considering the heterogeneous nature of home devices. We propose Halda to solve this NP-hard LDA problem, which assigns model layers to device CPUs and GPUs and minimizes token latency. On a small, heterogeneous, and budget-friendly home cluster, prima.cpp outperforms llama.cpp, exo, and dllama in speed, model size, and memory pressure. This work requires lower-end hardwares and has a better support for cross-platform deployment. With prima.cpp, we are able to collaborate phones, tablets, laptops, desktops etc. to bring 70B LLMs (Llama 3, DeepSeek R1, Qwen 2.5, QwQ) to home Siri.
Limitations: (a) Limited device types and quantity restrict our exploration of diverse home clusters. Readers are welcome to share their results on our Github repo. (b) Though much faster than other on-device systems, 70B models remain much slower than those on the cloud. Future work will integrate IQ1/Q4K for higher efficiency. (c) In low-RAM clusters without SSDs or GPUs, larger models will be extremely slow. (d) Token latency is heavily affected by memory competition. If there are other processes, prima.cpp slows down to free RAM for them and speeds up when they stop, so we use a stable value from multiple runs instead of an error bound. (e) Prima.cpp unlocks larger-scale open-source LLMs on user devices, where malicious content may not be filtered. The open-source community should enhance oversight of their models to prevent misuse.
@@ Appendix A Appendix
A.1 Prefetch-release Effect
@@
As illustrated in Figure 3. Before computation starts, the OS prefetches 3 model layers to the available memory limit. However, it does not stop and continues to load the 4th layer, causing the 1st layer to be released. This “prefetch-release" cycle repeats, so by the end, the last 3 layers are in memory, while the first 3 are not. Then, when computation begins, the 1st layer, which is not in memory, triggers a page fault, prompting the OS to reload it and the 4th layer to be released. Finally, all layers are loaded twice, adding unnecessary disk overhead without any benefit from prefetching.
@@
Figure 3: Illustration of model layers loaded into memory in pipeline parallelism with prefetching. In this case, the device handles 6 model layers but its available memory can only hold 3. The green blocks show the layers loaded into memory, while white blocks indicate those not yet loaded.
A.2 How Piped-ring Parallelism Solve the Prefetch-release Effect?
@@
To illustrate, Figure 4 considers a fast-disk device where prefetching is fast enough to complete before computation begins. In this case, with a fast disk and a layer window size of 2, ① prefetching is fast enough to load 2 layers before computation begins, then ② computation runs without page faults. Then, ③ the next round of 2 layers are prefetched, replacing the used ones. Steps ②-⑦ repeat until inference is complete. Prefetching overlaps with other devices’ operations, so its latency does not contribute to inference time. Here, with no page faults, total latency comes only from computation. In other words, disk loading latency is fully overlapped.
@@
Figure 4: Illustration of model layers loaded into memory in piped-ring parallelism with fast disk.
@@
Figure 5 shows a common case with a slow disk. In this case, ① prefetching loads only one layer then ② computation begins, ③ a page fault triggers when reaching the 2nd layer, blocking until it loads. After computation, ④ the device prefetches the next round of 2 layers, but only one layer loads due to the slow disk and releasing the oldest layer. Then, ⑤ the next round of computation begins, and ⑥ at the 6th layer, another page fault occurs. This cycle of “loading (prefetch) - computing - loading (page fault) - computing" repeats until inference completes. While page fault-induced loading blocks computation, prefetching helps overlap some latency.
@@
Figure 5: Illustration of model layers loaded into memory in piped-ring parallelism with slow disk.
@@
We use a timeline to visualize this overlap. In Figure 6, green blocks show prefetching that is overlapped, and orange blocks show page fault-induced loading that is not overlapped. In Figure 6(a), with a fast disk, disk loading is fully overlapped. In Figures 6(b) and 6(c), with a device has a slow disk, only part of the disk loading is overlapped, while others are fully overlapped. In Figures 6(c), 6(d) and 6(e), while disk loading is not fully hidden, piped-ring parallelism significantly reduces token latency compared to vanilla pipeline parallelism. In Figure 6(e), while prefetching is used, it exceeds memory limits and triggers “prefetch-release", where the OS releases the earlier prefetched layers as new ones load, adding disk cost with no benefit. This underscores the need to combine piped-ring parallelism with prefetching for higher efficiency.
@@
Figure 6: Timeline of (a,b) piped-ring parallelism on homogeneous devices with fast and slow disks; (c,e) piped-ring parallelism on heterogeneous devices with same and different window sizes; and (d,e) vanilla pipeline parallelism on heterogeneous devices with and without prefetching.
@@ A.3 Layer-to-Device Assignment: From Latency Analysis to Vectorized Model
Assume there are $M$ devices, where the layer window size for device $d_{m}$ is $w_{m}$. On device $d_{m}$, the number of GPU layers $n_{m}$ is defined as: within a layer window of size $w_{m}$, $n_{m}$ layers run on the GPU, while the remaining $w_{m}-n_{m}$ layers run on the CPU ($w_{m}$ and $n_{m}$ can vary across devices). Our objective is to find a vector ${\bm{w}}=\{w_{1},\cdots,w_{M}\}$ and a vector ${\bm{n}}=\{n_{1},\cdots,n_{M}\}$ to minimize the token latency $T$, which is the sum of latencies from computation $T_{m}^{\text{comp}}$, memory access $T_{m}^{\text{mem}}$, disk loading $T_{m}^{\text{disk}}$, and communication $T_{m}^{\text{comm}}$ on each device.
@@
Here, we minimize $T=\sum_{m=1}^{M}T_{m}$ instead of $T=\max\{T_{m}\}$ because Figure 6(f) is an idealized visualization. In practice, the OS does not start prefetching immediately after computation, and the timing is unknown. As a result, device 4 experiences more “bubbles" and higher page fault-induced latency than expected. This uncertainty prevents solving $T=\max\{T_{m}\}$ before deployment (it is also hard to measure), and historical data is useless due to fluctuating device conditions. Thus, we take a worst-case approach, assuming the OS hasn’t started prefetching when computation begins, leading to our objective $T=\sum_{m=1}^{M}T_{m}$. Next, we analyze these latencies in detail.
@@
Estimation of computation latency $T_{m}^{\text{comp}}$. The computation latency on device $d_{m}$ is defined as the time taken to process $l_{m}$ model layers and the output layer (if $d_{m}$ is the head device), where $l_{m}^{\text{gpu}}$ layers run on the GPU, and the remaining $l_{m}-l_{m}^{\text{gpu}}$ layers and output layer run on the CPU. Here, we have $l_{m}=\left\lfloor\frac{L}{W}\right\rfloor w_{m}+\min(w_{m},\max(0,R-\sum_{j=1}^{m-1}\min(w_{j},R)))$, $l_{m}^{\text{gpu}}=\left\lfloor\frac{L}{W}\right\rfloor n_{m}+\min(n_{m},\max(0,R-\sum_{j=1}^{m-1}\min(w_{j},R)))$, where $n_{m}\leq w_{m}$, $W=\sum_{m=1}^{M}w_{m}$, and $R=L\mod W$. Since the input layer adopts a look-up table, it does not contribute to computation latency.
@@
To estimate the computation time, we develop a model profiler to count the number of floating-point operations (FLOPs) for each model layer and a device profiler to measure the floating-point throughput (FLOPS) of each device. Taking Q4K as an example, the model weights are primarily quantized in the Q4K format, though some weights use other formats. Specifically, we consider $\mathcal{Q}=${Q4_K, Q5_K, Q6_K, Q8_0, F16, F32}, as well as three types of backends CPU, CUDA, and Metal. The FLOPs for each layer $\mathcal{F}_{m}$ and output layer $\mathcal{F}_{m}^{\text{out}}$ consists of 6 values $\mathcal{F}_{m}=\{f_{m}^{\text{q4k}},f_{m}^{\text{q5k}},f_{m}^{\text{q6k}},f_{m}^{\text{q80}},f_{m}^{\text{f16}},f_{m}^{\text{f32}}\}$, $\mathcal{F}_{m}^{\text{out}}=\{f_{m,\text{out}}^{\text{q4k}},f_{m,\text{out}}^{\text{q5k}},f_{m,\text{out}}^{\text{q6k}},f_{m,\text{out}}^{\text{q80}},f_{m,\text{out}}^{\text{f16}},f_{m,\text{out}}^{\text{f32}}\}$, with each represents the FLOPs under a specific quantization format. The FLOPS $\mathcal{S}_{m}$ consists of 3 sets $\{\mathcal{S}_{m}^{\text{cpu}},\mathcal{S}_{m}^{\text{cuda}},\mathcal{S}_{m}^{\text{metal}}\}$, with each set consists of 6 values (e.g., for CPU, $\mathcal{S}_{m}^{\text{cpu}}=\{s_{m}^{\text{cpu},\text{q4k}},s_{m}^{\text{cpu},\text{q5k}},s_{m}^{\text{cpu},\text{q6k}},s_{m}^{\text{cpu},\text{q80}},s_{m}^{\text{cpu},\text{f16}},s_{m}^{\text{cpu},\text{f32}}\}$) representing the floating-point throughput under specific device and quantization format. $\mathcal{F}_{m}$, $\mathcal{F}_{m}^{\text{out}}$ and $\mathcal{S}_{m}$ can be easily extended.
@@
$T_{m}^{\text{comp}}=(l_{m}-l_{m}^{\text{gpu}})\sum_{q\in\mathcal{Q}}\frac{f_{m}^{q}}{s_{m}^{\text{cpu},q}}+l_{m}^{\text{gpu}}\sum_{q\in\mathcal{Q}}\frac{f_{m}^{q}}{s_{m}^{\text{gpu},q}}+\mathbb{I}_{m=1}\cdot\sum_{q\in\mathcal{Q}}\frac{f_{m,\text{out}}^{q}}{s_{m}^{\text{cpu},q}},$ (12)
@@
Here, $s_{m}^{\text{gpu},q}$ refers to GPU FLOPS. If prima.cpp is compiled with CUDA support, $s_{m}^{\text{gpu},q}$ corresponds to $s_{m}^{\text{cuda},q}$. If it runs on an Apple device with Metal enabled, $s_{m}^{\text{gpu},q}$ corresponds to $s_{m}^{\text{metal},q}$. In our current implementation, the output layer is executed only on CPUs by the master node $m=1$.
@@
Estimation of memory access latency $T_{m}^{\text{mem}}$. This latency consists of three components: a) kvcache copy time $T_{m}^{\text{kv\_cpy}}$: the time taken to copy the new token cache to the key-value cache storage within a specific device; b) device copy time $T_{m}^{\text{dev\_cpy}}$: the time taken to copy hidden states between the CPU and the GPU (CUDA or Metal); c) device loading time $T_{m}^{\text{dev\_load}}$: the time taken to load data from RAM or VRAM into the processing cores of the CPU or GPU.
@@
For the kvcache copy time, in each token step, new key and value caches are generated with dimensions $(h_{k}e_{k},1)$ and $(h_{v}e_{v},1)$, respectively. Here, $h_{k}$ and $h_{v}$ are the number of attention heads for the key and value caches, $e_{k}$ and $e_{v}$ are the embedding size per head for the key and value vectors. Thus, for generating one token, each layer needs to copy $h_{k}e_{k}+h_{v}e_{v}$ values to the key-value cache storage. If the values are stored in the F16 format, each value takes 2 bytes, so the total number of bytes to be copied is $2(h_{k}e_{k}+h_{v}e_{v})$. In the device profiler module, we measure the time of copying $2(h_{k}e_{k}+h_{v}e_{v})$ bytes within CPU, CUDA, and Metal to obtain $t_{m}^{\text{kv\_cpy,cpu}}$ and $t_{m}^{\text{kv\_cpy,gpu}}$. Then, the kvcache copy time $T_{m}^{\text{kv\_cpy}}$ can be estimated by $(l_{m}-l_{m}^{\text{gpu}})t_{m}^{\text{kv\_cpy,cpu}}+l_{m}^{\text{gpu}}t_{m}^{\text{kv\_cpy,gpu}}$.
@@
For the device copy time, this latency arises when the GPU is enabled, as it involves copying the input from RAM to VRAM and then copying the output from VRAM back to RAM. Both the input and output have dimensions $(e,1)$, where $e$ represents the embedding size. These values are typically stored in the F32 format. In the device profiler module, we measured the latency for two operations: the time taken to copy $4e$ bytes of data from RAM to VRAM, denoted as $t_{m}^{\text{ram-vram}}$, and the time taken to copy $4e$ bytes of data from VRAM to RAM, denoted as $t_{m}^{\text{vram-ram}}$. For a sequence of layers within a window, one RAM-to-VRAM copy and one VRAM-to-RAM copy are needed, so the device copy time for one window is $t_{m}^{\text{ram-vram}}+t_{m}^{\text{vram-ram}}$. For device $d_{m}$, it was assigned $\mathcal{W}_{m}=\left\lfloor\frac{L}{W}\right\rfloor+\min(1,\max(0,R-\sum_{j=1}^{m-1}\min(w_{j},R)))$ windows. Thus, the device copy time for device $d_{m}$ is $T_{m}^{\text{dev\_cpy}}=\mathcal{W}_{m}(t_{m}^{\text{ram-vram}}+t_{m}^{\text{vram-ram}})(1-\mathbb{I}_{m}^{\text{UMA}})$, where $\mathbb{I}_{m}^{\text{UMA}}=1$ indicates that device $d_{m}$ uses a unified memory architecture (UMA, e.g., Apple with M1 chip) and the CPU and GPU use the shared memory, so no explicit RAM-VRAM copy is needed.

### v2

@@ 4.2 Ablation Study on Prefetching, Halda, and Pipelined-Ring Parallelism
8B 2.0% 20.0% 51.3% 42.5% 13.5% 12.8% 55.8% 12.8% 5.3% 5.4% 2.7% $\leq$1.0%
14B 2.5% - - - - - - - 5.3% 4.3% 2.2% $\leq$1.0%
30B 8.0% - - - - - - - 3.0% 5.7% 2.9% $\leq$1.0%
45B 3.9% - - - - - - - 4.9% $\leq$1.0% 6.0% $\leq$1.0%
60B 5.5% - - - - - - - 6.3% 4.7% 4.7% $\leq$1.0%
65B 15.6% - - - - - - - 3.9% $\leq$1.0% $\leq$1.0% $\leq$1.0%
70B 6.0% OOM OOM OOM OOM OOM OOM OOM 4.7% 4.8% 4.8% $\leq$1.0%
4.3 Low Memory Pressure to Preserve User Experience
Memory pressure is critical for user experience, as high pressure can slow apps or crash the device. For example, if a phone runs an LLM service while the user is browsing TikTok, memory competition may cause TikTok to lag or crash, prompting the user to terminate the LLM. An implicit advantage of prima.cpp is that it accounts for this by running at a lower memory priority than other apps: it reduces RAM usage when apps start and uses more when they stop. By giving up RAM to keep other apps responsive, prima.cpp preserves user experience and protects itself from being killed by the user.
We define memory pressure from the LLM as $\Delta\texttt{mem\_available}/\texttt{mem\_total}$ (e.g., 2 GiB used / 8 GiB total = 25%)111111mem_available includes free and reclaimable pages, so its reduction reflects only non-reclaimable pages, i.e., true memory stress. Appendix A.8 shows each device’s memory footprint, but not memory pressure, since reclaimable memory (e.g., page cache) is included and can be freed instantly.. Table 5 shows that exo and dllama pin model weights in RAM, raising memory pressure by over 50% on some devices. This can force the OS to reclaim memory from other apps (e.g., compression, swapping), causing lag or crashes, and can still lead to OOM for larger models. The root cause is that they prioritize the LLM at the cost of other apps, which users do not want. In contrast, llama.cpp and prima.cpp exhibit low pressure: both pin only minimal KV caches and compute buffers, while prima.cpp further distributes them across devices. Model weights are loaded via mmap and cached in pages that can be released instantly for other apps. These properties make prima.cpp a practical choice for consumer devices, where LLMs must run alongside user apps.
@@ 5 Conclusion
This paper proposes prima.cpp, the first on-device distributed system to deliver practical speed for 30-70B LLMs on consumer-grade home devices, achieving 26 tokens/s for 32B models and 2 tokens/s for 70B models with speculative decoding. We introduce PRP to run models exceeding the memory limit and resolve the prefetch-release conflict to reactivate prefetching. To handle device heterogeneity and reduce bubbles in PRP, we develop the LDA model, which captures the heterogeneity in computing, communication, memory, and OS-specific reclamation, disk optimizations. We solve it via Halda, which performs smart layer partitioning for minimal TPOT and automatically selects the best-performing devices. The system is meticulously designed, and this level of detail surpasses prior work that often relies on simplified heuristics. It offers features desired by home users: no extra hardware cost, no specialized hardware, privacy, offline, no queueing or timeouts, models beyond memory limit, practical intelligence and speed, low memory pressure, cross-platform, heterogeneity awareness, automatic workload allocation and device selection, and Wi-Fi ready. These make prima.cpp a compelling choice for home deployment and drive broader adoption of embodied AI at home.
@@ Appendix A Appendix
A.1 Prefetch-release Conflict
@@
As illustrated in Fig. 3, before computation starts, the OS prefetches 3 model layers to the available memory limit. However, it does not stop and continues to load the 4th layer, causing the 1st layer to be released. This prefetch-release cycle repeats, so by the end, the last 3 layers are in memory, while the first 3 are not. Then, when computation begins, the 1st layer, which is not in memory, triggers a page fault, prompting the OS to reload it and the 4th layer to be released. Finally, all layers are loaded twice, incurring unnecessary disk I/O without any benefit from prefetching.
@@
Figure 3: Illustration of model layers loaded into memory in PP with prefetching. In this case, the device handles 6 model layers, but its available memory can only hold 3. The green blocks show the layers loaded into memory, while white blocks indicate those not yet loaded.
A.2 How Pipelined-ring Parallelism Solves the Prefetch-release Conflict
@@
Fig. 4 illustrates a fast-disk device where prefetching is fast enough to complete before computation begins. In this case, with a fast disk and a layer window size of 2, ① prefetching is fast enough to load 2 layers before computation begins, then ② computation runs without page faults. Then, ③ the next round of 2 layers is prefetched, replacing the used layers. Steps ②-⑦ repeat until inference is complete. Prefetching overlaps with other devices’ operations, so its latency does not contribute to TPOT. Here, with no page faults, TPOT comes only from computation. In other words, disk loading latency is fully overlapped.
@@
Figure 4: Illustration of model layers loaded into memory in PRP with a fast disk.
@@
Fig. 5 shows a common case with a slow disk. In this case, ① prefetching loads only one layer, then ② computation begins, ③ a page fault is triggered upon reaching the 2nd layer, blocking until it loads. After computation, ④ the device prefetches the next round of 2 layers, but only one layer loads due to the slow disk, and the OS releases the oldest layer. Then, ⑤ the next round of computation begins, and ⑥ at the 6th layer, another page fault occurs. This cycle of "loading (prefetch) - computing - loading (page fault) - computing" repeats until inference completes. While page fault-induced loading blocks computation, prefetching helps overlap some latency.
@@
Figure 5: Illustration of model layers loaded into memory in PRP with a slow disk.
@@
We use a timeline to visualize this overlap. In Fig. 6, green blocks show prefetching that is overlapped, and orange blocks show page fault-induced loading that is not overlapped. In Fig. 6a, with a fast disk, disk loading is fully overlapped. In Figs. 6b and 6c, with a device that has a slow disk, only part of the disk loading is overlapped, while other devices are fully overlapped. In Figs. 6c, 6d, and 6e, although disk loading is not fully hidden, PRP significantly reduces TPOT compared with vanilla PP. In Fig. 6e, while prefetching is used, it exceeds memory limits and triggers prefetch-release, where the OS releases earlier prefetched layers as new ones are loaded, adding disk I/O cost without benefit. This underscores the need to combine PRP with prefetching for higher efficiency.
@@
Figure 6: Timeline of (a,b) PRP on homogeneous devices with fast/slow disks; (c,e) PRP on heterogeneous devices with the same/different window sizes; and (d,e) vanilla $P$ on heterogeneous devices with/without prefetching.
@@ A.3 Layer-to-Device Assignment: From Latency Analysis to Vectorized Model
Assume there are $M$ devices, where the layer window size for device $d_{m}$ is $w_{m}$. On device $d_{m}$, the number of GPU layers $n_{m}$ is defined as follows: within a layer window of size $w_{m}$, $n_{m}$ layers run on the GPU, while the remaining $w_{m}-n_{m}$ layers run on the CPU ($w_{m}$ and $n_{m}$ can vary across devices). Our objective is to find a vector ${\bm{w}}=\{w_{1},\cdots,w_{M}\}$ and a vector ${\bm{n}}=\{n_{1},\cdots,n_{M}\}$ to minimize the TPOT $T$, which is the sum of latencies from computation $T_{m}^{\text{comp}}$, memory access $T_{m}^{\text{mem}}$, disk loading $T_{m}^{\text{disk}}$, and communication $T_{m}^{\text{comm}}$ on each device.
@@
Here, we minimize $T=\sum_{m=1}^{M}T_{m}$ instead of $T=\max\{T_{m}\}$, because Fig. 6f is an idealized illustration. In practice, the OS does not start prefetching immediately after computation ends, and the timing is unknown. As a result, device 4 may experience more bubbles and higher page fault-induced latency than expected. This uncertainty prevents us from solving $T=\max\{T_{m}\}$ before deployment (which is also hard to measure), and historical data is useless due to fluctuating device conditions. Thus, we take a worst-case approach, assuming the OS has not started prefetching when computation begins, leading to our objective $T=\sum_{m=1}^{M}T_{m}$. Next, we analyze these latencies in detail.
@@
Estimation of computation latency $T_{m}^{\text{comp}}$. The computation latency on device $d_{m}$ is defined as the time taken to process $l_{m}$ model layers and the output layer (if $d_{m}$ is the head device), where $l_{m}^{\text{gpu}}$ layers run on the GPU, and the remaining $l_{m}-l_{m}^{\text{gpu}}$ layers and output layer run on the CPU. Here, we have $l_{m}=\left\lfloor\frac{L}{W}\right\rfloor w_{m}+\min(w_{m},\max(0,R-\sum_{j=1}^{m-1}\min(w_{j},R)))$, $l_{m}^{\text{gpu}}=\left\lfloor\frac{L}{W}\right\rfloor n_{m}+\min(n_{m},\max(0,R-\sum_{j=1}^{m-1}\min(w_{j},R)))$, where $n_{m}\leq w_{m}$, $W=\sum_{m=1}^{M}w_{m}$, and $R=L\mod W$. Since the input layer uses a lookup table, it does not contribute to computation latency.
@@
To estimate the computation time, we develop a model profiler to count the floating-point operations (FLOPs) for each model layer and a device profiler to measure the floating-point throughput (FLOPS) of each device. Taking Q4K as an example, the model weights are primarily quantized in the Q4K format, though some weights use other formats. Specifically, we consider $\mathcal{Q}=${Q4_K, Q5_K, Q6_K, Q8_0, FP16, FP32} and three types of backends: CPU, CUDA, and Metal. The FLOPs for each layer $\mathcal{F}_{m}$ and output layer $\mathcal{F}_{m}^{\text{out}}$ consist of 6 values: $\mathcal{F}_{m}=\{f_{m}^{\text{q4k}},f_{m}^{\text{q5k}},f_{m}^{\text{q6k}},f_{m}^{\text{q80}},f_{m}^{\text{fp16}},f_{m}^{\text{fp32}}\}$, $\mathcal{F}_{m}^{\text{out}}=\{f_{m,\text{out}}^{\text{q4k}},f_{m,\text{out}}^{\text{q5k}},f_{m,\text{out}}^{\text{q6k}},f_{m,\text{out}}^{\text{q80}},f_{m,\text{out}}^{\text{fp16}},f_{m,\text{out}}^{\text{fp32}}\}$, each representing the FLOPs under a specific quantization format. The FLOPS $\mathcal{S}_{m}$ consists of 3 sets: $\{\mathcal{S}_{m}^{\text{cpu}},\mathcal{S}_{m}^{\text{cuda}},\mathcal{S}_{m}^{\text{metal}}\}$, with each set consisting of 6 values (e.g., for CPU, $\mathcal{S}_{m}^{\text{cpu}}=\{s_{m}^{\text{cpu},\text{q4k}},s_{m}^{\text{cpu},\text{q5k}},s_{m}^{\text{cpu},\text{q6k}},s_{m}^{\text{cpu},\text{q80}},s_{m}^{\text{cpu},\text{fp16}},s_{m}^{\text{cpu},\text{fp32}}\}$) representing the floating-point throughput for a specific backend-quantization pair. $\mathcal{F}_{m}$, $\mathcal{F}_{m}^{\text{out}}$, and $\mathcal{S}_{m}$ can be easily extended for more backends and quantization formats.
@@
$T_{m}^{\text{comp}}=(l_{m}-l_{m}^{\text{gpu}})\sum_{q\in\mathcal{Q}}\frac{f_{m}^{q}}{s_{m}^{\text{cpu},q}}+l_{m}^{\text{gpu}}\sum_{q\in\mathcal{Q}}\frac{f_{m}^{q}}{s_{m}^{\text{gpu},q}}+\mathbb{I}_{m=1}\cdot\sum_{q\in\mathcal{Q}}\frac{f_{m,\text{out}}^{q}}{s_{m}^{\text{cpu},q}}.$ (12)
@@
Here, $s_{m}^{\text{gpu},q}$ refers to GPU FLOPS. If prima.cpp is compiled with CUDA support, $s_{m}^{\text{gpu},q}$ corresponds to $s_{m}^{\text{cuda},q}$. If it runs on an Apple device with Metal enabled, $s_{m}^{\text{gpu},q}$ corresponds to $s_{m}^{\text{metal},q}$. In our implementation, the output layer is executed on the CPU by the master node ($m=1$).
@@
Estimation of memory access latency $T_{m}^{\text{mem}}$. This latency consists of three components: (a) KV cache copy time $T_{m}^{\text{kv\_cpy}}$: the time taken to copy the new token’s cache to the KV cache storage on device $m$; (b) device copy time $T_{m}^{\text{dev\_cpy}}$: the time taken to copy hidden states between the CPU and the GPU (e.g., CUDA and Metal); (c) device loading time $T_{m}^{\text{dev\_load}}$: the time taken to load data from RAM or VRAM into the processing cores of the CPU or GPU.
@@
For the KV cache copy time, in each token step, new key and value caches are generated with dimensions $(h_{k}e_{k},1)$ and $(h_{v}e_{v},1)$, respectively. Here, $h_{k}$ and $h_{v}$ are the number of attention heads for the key and value caches, $e_{k}$ and $e_{v}$ are the embedding size per head for the key and value vectors, respectively. Thus, for generating one token, each layer needs to copy $h_{k}e_{k}+h_{v}e_{v}$ values to the KV cache storage. If values are stored in FP16, each takes 2 bytes, so the total number of bytes to be copied is $2(h_{k}e_{k}+h_{v}e_{v})$ bytes. In the device profiler module, we measure the time of copying $2(h_{k}e_{k}+h_{v}e_{v})$ bytes on CPU, CUDA, and Metal to obtain $t_{m}^{\text{kv\_cpy,cpu}}$ and $t_{m}^{\text{kv\_cpy,gpu}}$. Then, the KV cache copy time $T_{m}^{\text{kv\_cpy}}$ can be estimated by $(l_{m}-l_{m}^{\text{gpu}})t_{m}^{\text{kv\_cpy,cpu}}+l_{m}^{\text{gpu}}t_{m}^{\text{kv\_cpy,gpu}}$.
@@
For the device copy time, this latency arises when the GPU is enabled, as it involves copying the input from RAM to VRAM and then copying the output from VRAM back to RAM. Both input and output have shape $(e,1)$, where $e$ is the embedding size. These values are typically stored in the FP32 format. In the device profiler module, we measure the latency for two operations: the time taken to copy $4e$ bytes of data from RAM to VRAM, denoted as $t_{m}^{\text{ram-vram}}$, and the time taken to copy $4e$ bytes of data from VRAM to RAM, denoted as $t_{m}^{\text{vram-ram}}$. For a sequence of layers within a window, one RAM-to-VRAM copy and one VRAM-to-RAM copy are needed, so the device copy time for one window is $t_{m}^{\text{ram-vram}}+t_{m}^{\text{vram-ram}}$. For device $d_{m}$, it was assigned $\mathcal{W}_{m}=\left\lfloor\frac{L}{W}\right\rfloor+\min(1,\max(0,R-\sum_{j=1}^{m-1}\min(w_{j},R)))$ windows. Thus, the device copy time for $d_{m}$ is $T_{m}^{\text{dev\_cpy}}=\mathcal{W}_{m}(t_{m}^{\text{ram-vram}}+t_{m}^{\text{vram-ram}})(1-\mathbb{I}_{m}^{\text{UMA}})$, where $\mathbb{I}_{m}^{\text{UMA}}=1$ indicates that device $d_{m}$ uses a unified memory architecture (UMA, e.g., Apple M-series) and the CPU and GPU share memory, so no explicit RAM-VRAM copy is needed.